Shift report 20011004 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Oct 04 2001 - 00:13:24 EDT

  • Next message: Apache: "Shift report 20011004 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Oct  4 00:13:24 2001, Shift: 16:00-24:00
    Supervisor: burt budick, with: none
    Log book#9 pages 85 - 86
    
    Shift Summary:
    ------------------------------------------------------------
    Beam was on for first 2.5 hours
    
    Run 5111 was completed with 22 sequences, trigger 1 278K.
    
    Several short runs were devoted to a study the efficiency
     of T5 as a function of discriminator threshold.
    
    When the beam was dumped at 6:30 pm, pedestal runs for all
     detectors including TPC's were taken, as well as a drift
     velocity monitor run.
    
    Same problem for slat 3 in H2 as noted in previous run.
    
    Problem communicating with mainframe for H1 and C1, 
    although Supermon indicated both detectors functioning
     normally.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Oct 04 2001 - 01:12:54 EDT